OpenStudy (anonymous):

solve the exponential equation express the solution set in terms of natural logarithims 2^x+8=3

Nnesha (nnesha):

do you know the power rule of log(ln) ????

OpenStudy (anonymous):

\[2^{x+8}=3\] right ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

by the change o' base formula you get \[x+8=\frac{\ln(3)}{\ln(2)}\] and you can solve that for \(x\) in one step

OpenStudy (anonymous):

\[b^{\xi}=A\iff \xi =\frac{\ln(A)}{\ln(b)}\] is the change o' base

OpenStudy (anonymous):

ok thanks got it

OpenStudy (anonymous):

yw